OpenStudy (anonymous):

Solve the equation: t/3-1/2=1

OpenStudy (anonymous):

t/3 -1/2+(1/2) = 1 +(1/2)

OpenStudy (anonymous):

t/3 = 3/2

OpenStudy (anonymous):

(t/3)*3= (3/2)*3 t=9/2